# Multi-Derivative Runge-Kutta Flux Reconstruction for hyperbolic conservation laws

### v2

@@ 7 Numerical results
In this section, we test the MDRK scheme with numerical experiments using polynomial degree $N=3$ in all results. Most of the test cases from [3, 4] were tried and were seen to validate our claims; but we only show the important results here. We also tested all the benchmark problems for higher order methods in [20], and show some of the results from there.
@@ 7.3.2 Rotational low density problem
    and the near-field solution is defined for $r<2t\sqrt{p^{\prime}(\rho_{0})}$
@@
    $\operatorname{curl}(U,V)=V_{x}-U_{y}=-\frac{1}{2t}\neq 0$
    and the solution has one-point vacuum at the origin $r=0$. This is typical hurricane-like solution that behaves highly singular, particularly near the origin $r=0$. There are two issues here challenging the numerical schemes: one is the presence of the vacuum state which examines whether a high order scheme can keep the positivity preserving property; the other is the rotational velocity field for testing whether a numerical scheme can preserve the symmetry. In this regime, we take $v_{0}=10$ on the computational domain $[-1,1]^{2}$ with $\Delta x=\Delta y=1/100$. The boundary condition is given by the far field solution in (41).
@@ 7.3.4 Rayleigh-Taylor instability
$(\rho,u,v,p)=\begin{cases}(2,0,-0.025a\cos(8\pi x),2y+1),\quad&y\leq 0.5,\\ (1,0,-0.025a\cos(8\pi x),y+1.5),&y>0.5\end{cases}$
@@ 8 Summary and conclusions
This work introduces the fourth order multiderivative Runge-Kutta (MDRK) scheme of [18] in the conservative, quadrature free, Lax-Wendroff Flux Reconstruction framework to solve hyperbolic conservation laws. The idea is to cast each MDRK stage as an evolution involving a time average flux which is approximated by the Jacobian-free approximate Lax-Wendroff procedure. The D2 dissipation and EA flux of [3] have been introduced for numerical flux computation, which enhance stable CFL numbers and accuracy for nonlinear problems, respectively. The stable CFL numbers are computed using Fourier stability analysis for two commonly used correction functions, $g_{\operatorname{Radau}}$ and $g_{2}$, showing the improved CFL numbers. Numerical convergence studies for non-linear problems were performed, which revealed that optimal convergence rates were only seen when using the EA flux. The shock capturing blending scheme of [4] has also been introduced for the MDRK scheme applied at each stage. Following [4], flux limiting is applied to each MDRK stage to obtain admissibility preservation in means. The scaling limiter of [41] is then applied to obtain admissibility preservation at the solution points. Along with being provably admissibility preserving, the scheme is more capable of capturing small scale structures through the use of Gauss-Legendre solution points and application of MUSCL-Hancock scheme on the subcells. The claims are validated by numerical experiments for compressible Euler’s equations with the modern test suite [20] for high order methods.
@@ Appendix A Derivation of 2-stage, 4-th order scheme
$\partial_{t}\bm{L}\left(\bm{u}\right)=\bm{L}_{\bm{u}}\bm{u}_{t}$
@@
$\partial_{t}\bm{L}\left(\bm{u}^{\ast}\right)=\bm{L}_{\bm{u}}\left(\bm{u}^{\ast}\right)\bm{L}\left(\bm{u}^{\ast}\right)$
@@ Appendix B Treatment of source terms
$\bm{u}_{t}+\bm{f}\left(\bm{u}\right)_{x}=\text{{{s}}}$ (60)
where $\text{{{s}}}=\text{{{s}}}(\bm{u},t,\bm{x})$. The idea is to apply the MDRK scheme of [18] to (60) as follows
@@
$\begin{split}\mathbf{S}&:=\text{{{s}}}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\text{{{s}}}\left(\bm{u}^{n}\right)\approx\frac{1}{\Delta t/2}\int_{t^{n}}^{t^{n+\frac{1}{2}}}\text{{{s}}}\\ \mathbf{S}^{*}&:=\text{{{s}}}(\bm{u}^{n})+\frac{1}{6}\Delta t\left(\frac{\partial}{\partial t}\text{{{s}}}(\bm{u}^{n})+2\frac{\partial}{\partial t}\text{{{s}}}(\bm{u}^{\ast})\right)\approx\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\text{{{s}}}\end{split}$
@@
$\mathsf{u}^{(1)}$ $=-\frac{\Delta t}{\Delta x_{e}}\mathsf{D}\mathsf{f}+\Delta t\text{{{s}}}$
@@
${\mathsf{s}}^{(1)}$ $=\frac{1}{12}\left[-\text{{{s}}}(\mathsf{u}+2\mathsf{u}^{(1)})+8\text{{{s}}}(\mathsf{u}+\mathsf{u}^{(1)})-8\text{{{s}}}(\mathsf{u}-\mathsf{u}^{(1)})+\text{{{s}}}(\mathsf{u}-2\mathsf{u}^{(1)})\right]$
@@
${\mathsf{s}}^{*(1)}$ $=\frac{1}{12}\left[-\text{{{s}}}(\mathsf{u}^{*}+2\mathsf{u}^{*(1)})+8\text{{{s}}}(\mathsf{u}^{*}+\mathsf{u}^{*(1)})-8\text{{{s}}}(\mathsf{u}^{*}-\mathsf{u}^{*(1)})+\text{{{s}}}(\mathsf{u}^{*}-2\mathsf{u}^{*(1)})\right]$
@@ Acknowledgments
The work of Arpit Babbar and Praveen Chandrashekar is supported by the Department of Atomic Energy, Government of India, under project no. 12-R&D- TFR-5.01-0520.

### v3

@@ 8 Numerical results
In this section, we test the MDRK scheme with numerical experiments using polynomial degree $N=3$ in all results. Most of the test cases from [5, 6] were tried and were seen to validate our claims; but we only show the important results here. We also tested all the benchmark problems for higher order methods in [23], and show some of the results from there. The developments have been contributed to the Julia package Tenkai.jl [4] which has been used for all the numerical experiments.
@@ 8.3.2 Rotational low density problem
    and the near-field solution is defined for $r<2t\sqrt{p^{\prime}(\rho_{0})}$ as
@@
    $\text{curl}(U,V)=V_{x}-U_{y}=-\frac{1}{2t}\neq 0$
    and the solution has one-point vacuum at the origin $r=0$. This is a typical hurricane-like solution that behaves highly singular, particularly near the origin $r=0$. There are two issues here challenging the numerical schemes: one is the presence of the vacuum state which examines whether a high order scheme can keep the positivity preserving property; the other is the rotational velocity field for testing whether a numerical scheme can preserve the symmetry. In this regime, we take $v_{0}=10$ on the computational domain $[-1,1]^{2}$ with $\Delta x=\Delta y=1/100$. The boundary condition is given by the far field solution in (45).
@@ 8.3.4 Rayleigh-Taylor instability
$(\rho,u,v,p)=\begin{cases}(2,0,-0.025a\cos(8\pi x),2y+1),\quad&y\leq 0.5\\ (1,0,-0.025a\cos(8\pi x),y+1.5),&y>0.5\end{cases}$
@@
8.3.5 Timing studies of the MDRK method
We will compare the proposed fourth order multiderivative Runge-Kutta method in Flux Reconstruction (MDRK-FR) framework with the standard Runge-Kutta Flux Reconstruction (RK-FR) using the five stage, fourth order Strong Stability Preserving Runge-Kutta (SSPRK) time discretization together with the blending limiter. The comparison is made using Gauss-Legendre solution points, with each scheme using its optimal CFL number in (44). The CFL number of the MDRK scheme is taken from Table 1. For the fourth order SSPRK-FR scheme, we performed a Fourier stability analysis to determine its optimal CFL number to be 0.215. The MDRK scheme uses the MUSCL-Hancock blending scheme, while the SSPRK scheme uses a blending limiter with second order MUSCL reconstruction on the subcells at each stage. The comparison has been made in Table 2 using our code Tenkai.jl [4] for all the Cartesian tests in Section 8.3. Table 2 is not meant to be a comparison between different hardware vendors, especially since the processors compared are not of the same generation. These measurements are only to show the performance comparison of MDRK and SSPRK across different architectures. While there is a spread in the timings based on the hardware, we see that MDRK is about 20% to 40% faster than Runge-Kutta scheme for most problems. A more realistic comparison should be done with MPI versions of these codes, but it is not part of the current work.
Test Intel Xeon Gold 5320 AMD Threadripper Apple M3 Pro
SSPRK MDRK Ratio SSPRK MDRK Ratio SSPRK MDRK Ratio
Double Mach 14902 10857 1.37 9617 7690 1.25 6737 4907 1.37
Rotational flow 1367 1102 1.24 971 793 1.22 705 533 1.32
Riemann problem 12095 11319 1.06 8662 8384 1.03 6232 5307 1.17
Rayleigh Taylor 9811 7989 1.22 7091 5271 1.34 4753 3368 1.41
Table 2: Wall Clock Time (WCT) in seconds of the numerical tests of 2-D Euler’s equations for multidervative Runge-Kutta (MDRK) method and Strong Stability Preserving Runge-Kutta (SSPRK) methods in flux reconstruction framework.
8.4 Extension to curvilinear grids
The extension to curvilinear meshes to the MDRK scheme is made analogously to [3]. The time step size is computed using a CFL based formula (88) from [3]. A more detailed description and an introduction of error based time stepping [3] for the MDRK scheme will be part of a future work.
8.4.1 Isentropic vortex test
This is a test with exact solution taken from [17] where the domain is specified by the following transformation from $[0,1]^{2}\to\Omega$
$\bm{x}(\xi,\eta)=\left(\begin{array}[]{c}\xi L_{x}-A_{x}L_{y}\sin(2\pi\eta)\\ \eta L_{y}+A_{y}L_{x}\sin(2\pi\xi)\end{array}\right)$
which is a distortion of the square $[0,L_{x}]\times[0,L_{y}]$ with sine waves of amplitudes $A_{x},A_{y}$. Following [17], we choose length $L_{x}=L_{y}=0.1$ and amplitudes $A_{x}=A_{y}=0.1$. The boundaries are set to be periodic. A vortex with radius $R_{v}=0.005$ is initialized in the curved domain with center $(x_{v},y_{v})=(L_{x}/2,L_{y}/2)$. The gas constant is taken to be $R_{\operatorname{gas}}=287.15$ and specific heat ratio $\gamma=1.4$ as before. The free stream state is defined by the Mach number $M_{0}=0.5$, temperature $T_{0}=300$, pressure $p_{0}=10^{5}$, velocity $u_{0}=M_{0}\sqrt{\gamma R_{\operatorname{gas}}T_{0}}$ and density $\rho_{0}=\frac{p_{0}}{R_{\operatorname{gas}}T_{0}}$. The initial condition $\bm{u}_{0}$ is given by
$\begin{gathered}(\rho,u,v,p)=\left(\rho_{0}\left(\frac{T}{T_{0}}\right)^{\frac{1}{\gamma-1}},u_{0}\left(1-\beta\frac{y-y_{v}}{R_{v}}e^{\frac{-r^{2}}{2}}\right),u_{0}\beta\frac{x-x_{v}}{R_{v}}e^{\frac{-r^{2}}{2}},\rho(x,y)R_{\operatorname{gas}}T\right)\\ T(x,y)=T_{0}-\frac{(u_{0}\beta)^{2}}{2C_{p}}e^{-r^{2}},\qquad r=\sqrt{(x-x_{v})^{2}+(y-y_{v})^{2}}/R_{v}\end{gathered}$
where $C_{p}=R_{\operatorname{gas}}\gamma/(\gamma-1)$ is the heat capacity at constant pressure and $\beta=0.2$ is the vortex strength. The vortex moves in the positive $x$ direction with speed $u_{0}$ so that the exact solution at time $t$ is $\bm{u}(x,y,t)=\bm{u}_{0}(x-u_{0}t,y)$ where $\bm{u}_{0}$ is extended outside $\Omega$ by periodicity. We simulate the propagation of the vortex for one time period $t_{p}=L_{x}/u_{0}$ and perform numerical convergence analysis for degree $N=3$ in Figure 16b, showing optimal rates in grid size versus $L^{2}$ error norm for all the conserved variables.
Refer to caption Refer to caption
(a) (b)
Figure 16: Convergence analysis for isentropic vortex problem with polynomial degree $N=3$. (a) Density plot, (b) $L^{2}$ error norm of conserved variables
8.4.2 Supersonic flow over cylinder
Supersonic flow over a cylinder is computed at a free stream Mach number of $3$ with the initial condition
$(\rho,u,v,p)=(1.4,3,0,1)$
Solid wall boundary conditions are used on the cylinder and at the top and bottom boundaries. A bow shock forms ahead of the cylinder which reflects across the solid walls and interacts with the small vortices forming in the wake of the cylinder. The setup of Löhner’s smoothness indicator [21] is taken from an example of Trixi.jl [26]
[…]
Figure 17: Mach 3 flow over cylinder using solution and mesh polynomial degree $N=4$ at $t=5$ (a) Initial mesh, (b) adaptively refined mesh at final time, (c) density plot at final time
@@ 9 Summary and conclusions
This work introduces the fourth order multiderivative Runge-Kutta (MDRK) scheme of [20] in the conservative, quadrature free, Flux Reconstruction framework to solve hyperbolic conservation laws. The idea is to cast each MDRK stage as an evolution involving a time average flux which is approximated by the Jacobian-free approximate Lax-Wendroff procedure. The D2 dissipation and EA flux of [5] have been introduced for numerical flux computation, which enhance stable CFL numbers and accuracy for nonlinear problems, respectively. The stable CFL numbers are computed using Fourier stability analysis for two commonly used correction functions, $g_{\text{Radau}}$ and $g_{2}$, showing the improved CFL numbers. Numerical convergence studies for nonlinear problems were performed, which revealed that optimal convergence rates were only seen when using the EA flux. The shock capturing blending scheme of [6] has also been introduced for the MDRK scheme applied at each stage. Following [6], flux limiting is applied to each MDRK stage to obtain admissibility preservation in means. The scaling limiter of [46] is then applied to obtain admissibility preservation at the solution points. Along with being provably admissibility preserving, the scheme is more capable of capturing small scale structures through the use of Gauss-Legendre solution points and application of the MUSCL-Hancock scheme on the subcells. The claims are validated by numerical experiments for compressible Euler’s equations with the modern test suite [23] for high order methods.
@@ Appendix A Derivation of 2-stage, 4-th order scheme
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{L}_{t}}\left(\bm{u}\right)=\bm{L}_{\bm{u}}\bm{u}_{t},$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{L}_{t}}\left(\bm{u}^{\ast}\right)=\bm{L}_{\bm{u}}\left(\bm{u}^{\ast}\right)\bm{L}\left(\bm{u}^{\ast}\right)$
@@ Appendix B Treatment of source terms
$\bm{u}_{t}+\bm{f}\left(\bm{u}\right)_{x}=\bm{s}$ (64)
where $\bm{s}=\bm{s}(\bm{u},t,\bm{x})$. The idea is to apply the MDRK scheme of [20] to (64) as follows
@@
$\begin{split}\mathbf{S}&:=\bm{s}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\bm{s}\left(\bm{u}^{n}\right)\approx\frac{1}{\Delta t/2}\int_{t^{n}}^{t^{n+\frac{1}{2}}}\bm{s}\\ \mathbf{S}^{*}&:=\bm{s}(\bm{u}^{n})+\frac{1}{6}\Delta t\left(\frac{\partial}{\partial t}\bm{s}(\bm{u}^{n})+2\frac{\partial}{\partial t}\bm{s}(\bm{u}^{\ast})\right)\approx\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\bm{s}\end{split}$
@@
$\mathsf{u}^{(1)}$ $=-\frac{\Delta t}{\Delta x_{e}}\mathsf{D}\mathsf{f}+\Delta t\bm{s}$
@@
${\mathsf{s}}^{(1)}$ $=\frac{1}{12}\left[-\bm{s}(\mathsf{u}+2\mathsf{u}^{(1)})+8\bm{s}(\mathsf{u}+\mathsf{u}^{(1)})-8\bm{s}(\mathsf{u}-\mathsf{u}^{(1)})+\bm{s}(\mathsf{u}-2\mathsf{u}^{(1)})\right]$
@@
${\mathsf{s}}^{*(1)}$ $=\frac{1}{12}\left[-\bm{s}(\mathsf{u}^{*}+2\mathsf{u}^{*(1)})+8\bm{s}(\mathsf{u}^{*}+\mathsf{u}^{*(1)})-8\bm{s}(\mathsf{u}^{*}-\mathsf{u}^{*(1)})+\bm{s}(\mathsf{u}^{*}-2\mathsf{u}^{*(1)})\right]$
@@ Acknowledgments
The work of Arpit Babbar and Praveen Chandrashekar is supported by the Department of Atomic Energy, Government of India, under project no. 12-R&D- TFR-5.01-0520. We thank Jalil Khan for benchmarking our method on the AMD processor.